[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Sep 29 09:20:51 UTC 2017


HSU Shift Log for Fri Sep 29 2017 05:20:44 GMT-0400 (EDT)

Shift: A

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: Got red banners back after submitting guidance for Maria, Lee,  
and AL99. The trackers seemed to come back OK, although I did not receive  
SHIPS/LGEM. Also, the SDM reported that they did not receive the inputs for  
the HWRF/HMON submissions. Called Monica to look into the problems. It was  
eventually determined that the compute file timestamp needed to be updated  
and the file resubmitted to fix both problems. However, not sure why the  
timestamp didn't get updated appropriately on the first try.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170929/bd6f0051/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list